# Beyond Empathy: Integrating Diagnostic and Therapeutic Reasoning with Large Language Models for Mental Health Counseling

### v1

@@ 4.3 Performance Comparison
To evaluate PsyLLM’s effectiveness against state-of-the-art models, we conducted a turn-based dialogue evaluation on the OpenR1-Psy test set. We compared PsyLLM against baselines including ChatCounselor Liu et al. (2023b), CPsyCounX Zhang et al. (2024a), DeepSeek-V3 DeepSeek-AI et al. (2024), and GPT-4o OpenAI (2023). Evaluation utilized the four key metrics: Comprehensiveness, Professionalism, Authenticity, and Safety (with max scores outlined in Table 3). This process was automated using Gemini-2.0-Flash, guided by detailed prompts (refer to Fig. 11 in the Appendix). Table 4 shows the results. PsyLLM consistently demonstrates strong performance across all dimensions, achieving the highest scores in Comprehensiveness (0.904), Authenticity (1.440), and Safety (0.945). Although GPT-4o slightly edges out in Professionalism (2.218), PsyLLM’s score (2.210) is very close. Overall, PsyLLM attains the highest Normalized Average score (0.607), surpassing all baselines, including GPT-4o (0.552) and DeepSeek-V3 (0.540). These findings indicate that fine-tuning on the OpenR1-Psy dataset effectively enhances a language model’s capacity to generate clinically relevant and therapeutic counseling responses.
@@
To validate the automatic evaluation, we conducted a human evaluation. Five psychology graduate students assessed a randomly selected subset of dialogue turns from the test set, rating responses from PsyLLM and the baseline models (ChatCounselor Liu et al. (2023b), CPsyCounX Zhang et al. (2024a), DeepSeek-V3 DeepSeek-AI et al. (2024), GPT-4o Hurst et al. (2024)). Responses were evaluated blindly based on the four dimensions: Comprehensiveness, Professionalism, Authenticity, and Safety. As Table 5 shows, human evaluation confirms that PsyLLM consistently outperforms all baselines, achieving the highest scores across all four dimensions and the Normalized Average (0.601). It demonstrates PsyLLM’s superior ability to generate human-preferred and clinically counseling responses.
@@ 4.4 Analysis
| Method | Metrics | ||||
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
| Prompt w/o Clinical Frame | 1.075 | 2.994 | 2.079 | 0.979 | 0.740 |
| Prompt w/o Therapy Guidance | 0.803 | 2.871 | 2.053 | 0.985 | 0.697 |
| OpenR1-Psy | 1.083 | 3.277 | 2.299 | 0.988 | 0.778 |
| Human | 2.879 | 2.890 | 2.990 | 2.672 | 3.000 | 0.962 |
| Automatic Evaluation | 2.806 | 2.710 | 2.660 | 2.970 | 2.882 | 0.935 |
@@
To understand how the size of the underlying language model affects counseling performance after fine-tuning on OpenR1-Psy, we trained PsyLLM variants using different scales of the Qwen3 model series: Qwen3-1.7B, Qwen3-3B, and Qwen3-8B Yang et al. (2025). These models were fine-tuned using the same methodology and dataset as the main PsyLLM model. As shown in Figure 4 (Left), there is a clear trend of improving performance with increasing model size. The Qwen3-8B model achieves the highest overall Normalized Average score (0.757), indicating superior performance compared to the smaller models. The Qwen3-3B model performs better than the 1.7B version, achieving a Normalized Average of 0.738, while the smallest Qwen3-1.7B consistently shows the lowest overall performance with a Normalized Average of 0.675. These results indicate that larger base models, even when fine-tuned on the same dataset, possess stronger inherent capabilities that translate into better overall performance in generating nuanced, professional, and authentic psychological counseling responses.
@@
To investigate the influence of training data volume, we trained PsyLLM using varying proportions (20%, 40%, 60%, 80%, and 100%) of the OpenR1-Psy dataset, keeping the base model (Qwen3-8B) consistent. As shown in Figure 4 (Right), the impact of data scale on the Normalized Average is not strictly monotonic. Performance generally improves from 20% (0.718) up to 80% (0.746), with 80% achieving the highest Normalized Average. However, the Normalized Average drops significantly when using the full 100% of the training data (0.656). The Safety metric remained consistently high across all data scales (around 0.99). This indicates that simply using more data for model training does not always yield the best performance.
Effect of Diagnostic Standard & Therapeutic Strategy.
We conducted an ablation study on data generation by omitting either diagnostic guidance (“Prompt w/o Clinical Frame”) or therapeutic guidance (“Prompt w/o Therapy Guidance”), fine-tuning PsyLLM on these ablated datasets and the full OpenR1-Psy data. As seen in Table 6, training on the full OpenR1-Psy dataset yields the best Normalized Average (0.778). Removing therapeutic guidance results in a notable performance drop (0.697), especially in Comprehensiveness and Professionalism. Removing diagnostic guidance also reduces performance (0.740), though less significantly. It confirms that diagnostic standards and therapeutic strategies are critical.
@@ 5 Conclusion
This paper addresses the critical limitations of existing LLMs for mental health support, specifically, their lack of explicit diagnostic reasoning (DSM/ICD aligned) and diverse therapeutic modality integration. We introduce PsyLLM, the first model designed to systematically integrate both, facilitated by a novel automated data synthesis pipeline creating the OpenR1-Psy dataset. This dataset incorporates explicit reasoning traces guided by diagnostic standards and multiple therapeutic frameworks (e.g., CBT, ACT, psychodynamic), validated through multi-dimensional filtering. Our new benchmark, assessing counseling quality across comprehensiveness, professionalism, authenticity, and safety, demonstrates that PsyLLM significantly outperforms state-of-the-art baselines.
Limitations
PsyLLM performs well on short to medium-length dialogues but may struggle with long-term context tracking, complex emotional transitions, or maintaining nuanced therapeutic pacing across extended multi-turn interactions. Furthermore, due to limited computational resources and the lack of access to high-memory GPUs, PsyLLM is currently based on a 8B-sized model. Larger models with greater capacity for reasoning and emotional understanding have not been explored in this study.
Ethics Statements
@@
Appendix A Related Work
A.1 Mental Support Chatbots
A.2 Deep Reasoning in LLMs
Deep reasoning in LLMs has attracted increasing attention, particularly since the introduction of CoT prompting Wei et al. (2022), which encourages models to generate intermediate reasoning steps before producing final answers. This technique has shown notable improvements on arithmetic and symbolic tasks, revealing the latent ability of LLMs to perform multi-step inference. Recent work has extended deep reasoning capabilities beyond symbolic computation into more complex domains such as math reasoning, medical understanding, multimodal reasoning, and emotional understanding Zhou et al. (2025b, a); Hu et al. (2025); Zhou et al. (2024). For example, Dou et al. (2025) proposes a dual-system framework inspired by Kahneman’s theory of fast and slow thinking, integrating reinforcement learning and vision-language reasoning to support adaptive decision-making. Similarly, reinforcement learning has been leveraged to enhance both text-based reasoning and multimodal emotion recognition, with modality-specific contributions disentangled through verifiable rewards Zhao et al. (2025). In mathematical contexts, Wang et al. (2025) introduces a model that bridges vision and code by aligning image understanding with code generation, achieving superior performance in structured geometry problems compared to models like GPT-4o. As research pushes toward more interactive and human-centric applications, deep reasoning in dialogue systems has emerged as a challenging yet crucial frontier. Maintaining coherence, adaptability, and emotional sensitivity across multiple conversational turns introduces new demands. To address this, Shu et al. (2025) reformulates CoT prompting as a dialogue between reasoning agents and applies reinforcement learning to improve both coherence and solution diversity. In the context of empathetic conversations, Xie et al. (2025) proposes an LPE framework that combines CoT with emotional perception from spoken input, thereby guiding emotionally intelligent response generation. Moreover, psychological structures have been introduced to support deep reasoning in personality-related tasks: Yang et al. (2023) models psychological questionnaires as structured CoT chains for personality trait detection through multi-turn dialogue, while Chen et al. (2025) introduces a graph-based reasoning visualization system to facilitate collaboration between AI and therapists in therapeutic settings.
| Topic | Model | Comprehensiveness | Professionalism | Authenticity | Safety |
| Self-growth | ChatCounselor | 0.000 | 0.774 | 0.226 | 0.419 |
@@ A.2 Deep Reasoning in LLMs
| Emotion & Stress | ChatCounselor | 0.016 | 0.873 | 0.337 | 0.504 |
@@
| Education | ChatCounselor | 0.000 | 0.794 | 0.305 | 0.412 |
@@
| Love & Marriage | ChatCounselor | 0.018 | 0.788 | 0.343 | 0.472 |
@@
| Family Relationship | ChatCounselor | 0.021 | 0.785 | 0.343 | 0.453 |
@@
| Social Relationship | ChatCounselor | 0.017 | 0.922 | 0.353 | 0.457 |
@@
| Sex | ChatCounselor | 0.000 | 0.667 | 0.333 | 0.667 |
| CPsyCounX | 0.000 | 0.833 | 0.333 | 0.500 | |
| DeepSeek-V3 | 0.833 | 2.167 | 1.167 | 1.000 | |
| GPT-4o | 0.667 | 2.000 | 0.667 | 1.000 | |
| PsyLLM | 0.500 | 2.000 | 1.500 | 0.667 | |
| Career | ChatCounselor | 0.000 | 0.837 | 0.279 | 0.395 |
@@
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
@@
| Comprehensiveness | Professionalism | Authenticity | Safety | Normalized Avg | |
@@
| Perspective | Criterion | Score | Subtotal |
| Comprehensiveness | 1.1 Accurate Empathy for Core Emotions: Does the response accurately identify and validate the client’s specific core emotions (e.g., anxiety, powerlessness, self-doubt) with precise language, going beyond generic phrases? | 1 | |
| 1.2 Insight into Underlying Needs & Pain Points: Does the response show insight into the client’s deeper, often unstated, emotional pain points or unmet needs (e.g., for safety, belonging, understanding) and gently guide exploration or affirm their validity? | 1 | 2 | |
| Professionalism | 2.1 Emotional Attunement & Validation: Does the counselor deeply empathize with and validate the client’s emotional experience, making the client feel "felt" and understood? | 1 | |
| 2.2 Natural Integration of Support: Is supportive guidance (reflection, affirmation, gentle questions) woven naturally into the conversation, without jargon or explicit theory-naming? | 1 | ||
| 2.3 Respectful & Non-Directive Style: Is the language consistently warm, respectful, and non-judgmental, avoiding commands or prescriptive "you should" statements? | 1 | 4 | |
| 2.4 Gentle Guidance & Client Agency: Are suggestions offered tentatively, encouraging client-led exploration and small steps, rather than providing direct solutions and respecting client autonomy? | 1 | ||
| Authenticity | 3.1 Conversational Fluency: Does the counselor communicate in a fluent, natural way, mirroring real therapeutic dialogue? | 1 | |
| 3.2 Relational Presence: Is there evidence of genuine emotional connection and relational presence with the client? | 1 | 3 | |
| 3.3 Pacing and Responsiveness: Does the counselor maintain appropriate pacing, leaving space for client expression, and respond in a natural, conversational manner to the client’s immediate input? | 1 | ||
| Safety | 4.1 Ethical Boundaries & Client Protection: Does the counselor avoid intrusive or leading language, respect privacy by not probing for unnecessary sensitive details, and prioritize client well-being and autonomy? | 0.5 | |
| 4.2 Fostering Psychological Safety: Does the counselor consistently convey unconditional respect and acceptance, creating an environment free of judgment or dismissal, where the client feels safe to be authentic? | 0.5 | 1 |

### v2

@@ 4.3 Performance Comparison
To evaluate PsyLLM’s effectiveness against state-of-the-art models, we conducted a turn-based dialogue evaluation on the OpenR1-Psy test set. We compared PsyLLM against baselines including ChatCounselor Liu et al. (2023b), CPsyCounX Zhang et al. (2024a), DeepSeek-V3 DeepSeek-AI et al. (2024), PsyDTLLM Xie et al. (2025a), MeChat Qiu et al. (2024a) and GPT-4o OpenAI (2023). Evaluation utilized the four key metrics: Emp&In, Sup&Aut, Att&Pre, and Saf&Bou (with their respective maximum scores detailed in the caption of Table 4).
Automatic Evaluation.
This process was automated using Gemini-2.0-Flash, guided by detailed prompts. To validate the reliability and effectiveness of our LLM-based evaluation, we calculate the Spearman rank correlation coefficient Sedgwick (2014) with ratings from psychology experts. The overall assessment (OA) shows a significant positive correlation of 0.605, demonstrating strong agreement between automatic and expert evaluations (Further details are provided in Appendix C).
As shown in Table 4, PsyLLM consistently demonstrates strong performance across all dimensions, achieving the highest scores in Emp&In (0.904), Att&Pre (1.440), and Saf&Bou (0.945). Although GPT-4o slightly edges out in Sup&Aut (2.218), PsyLLM’s score (2.210) is very close. Overall, PsyLLM attains the highest Normalized Average score (0.607), surpassing all baselines, including GPT-4o (0.552) and DeepSeek-V3 (0.540). Figure 4 further shows detailed scores of PsyLLM and other baselines. Across the topic distribution, PsyLLM consistently outperforms competing methods. This highlights its practical utility for psychological counseling.
@@
To validate the automatic evaluation, we conducted a human evaluation. Five psychology experts with clinical training evaluated a randomly selected subset of 100 dialogue turns per model from the test set, rating the responses generated by PsyLLM and the baseline models. Responses were evaluated blindly based on our four evaluation dimensions. As Table 4 shows, human evaluation confirms that PsyLLM consistently outperforms all baselines, achieving the highest scores across all four dimensions and the Normalized Average (0.601). It shows PsyLLM’s superior ability to generate human-preferred and clinically counseling responses.
| Automatic | 2.806 | 2.710 | 2.660 | 2.970 | 2.882 | 0.935 |
| Human | 2.879 | 2.890 | 2.990 | 2.672 | 3.000 | 0.962 |
Comparison of Reasoning Integration Strategies.
To validate our joint reasoning-response generation methodology, we compared PsyLLM with two alternative strategies on the OpenR1-Psy test set. The baselines include: (1) In-Context Learning (ICL), which provides few-shot demonstrations from our dataset to the base model without fine-tuning; and (2) a Two-Phase Prompt method, where the model first generates a reasoning trace and then, in a separate step, generates the counselor response based on the input and the trace.
As shown in Figure 5, our method (PsyLLM) significantly outperforms both ICL and the Two-Phase Prompt across all evaluation metrics. The results highlight that Supervised Fine-Tuning is essential for embedding complex therapeutic reasoning, a task where ICL is insufficient. More importantly, the superiority over the two-phase approach confirms the benefit of our joint generation strategy. By training the model to produce the reasoning () and response () in a single, continuous output, we enforce a tight coupling between clinical rationale and the final intervention. It ensures the response is directly grounded in its preceding reasoning, better emulating a fluid and coherent human clinical thought process.
@@ 4.4 Analysis
Effect of Diagnostic Standard & Therapeutic Strategy.
We conducted an ablation study on data generation by omitting either diagnostic guidance (“Prompt w/o Clinical Frame”) or therapeutic guidance (“Prompt w/o Therapy Guidance”), fine-tuning PsyLLM on these ablated datasets and the full OpenR1-Psy data. Figure 6 (Left) shows that using the full OpenR1-Psy dataset yields the best performance (0.778). Excluding therapeutic guidance leads to a notable drop (0.697), especially in Sup&Aut and Emp&In; removing diagnostic guidance also reduces performance (0.740). It confirms that diagnostic standards and therapeutic strategies are critical.
Correlation Analysis of Evaluation Metrics.
To better understand the relationships between our evaluation dimensions, we conducted a correlation analysis of the metrics. The correlation matrix, depicted in Figure 6 (Right), illustrates the interdependencies among Empathy & Insight (Com), Support & Autonomy (Pro), Attunement & Presence (Aut), and Safety & Boundaries (Saf). The analysis reveals strong positive correlations between several core therapeutic qualities. Specifically, Support & Autonomy shows a high correlation with Attunement & Presence (0.73), and Empathy & Insight is strongly correlated with Support & Autonomy (0.70). This suggests that counseling responses that are well-attuned and demonstrate presence are also highly likely to be perceived as supportive and autonomy-fostering. Similarly, responses that convey deep empathy and insight are also effective in providing support. In contrast, the Safety & Boundaries metric exhibits significantly lower correlations with the other dimensions (0.33 with Empathy & Insight, 0.25 with Support & Autonomy, and 0.20 with Attunement & Presence). This indicates that safety is a distinct and foundational component of a counseling response that does not automatically arise from qualities like empathy or support.
@@
To understand how the size of the underlying language model affects counseling performance after fine-tuning on OpenR1-Psy, we trained PsyLLM variants using different scales of the Qwen3 model series: Qwen3-1.7B, Qwen3-3B, and Qwen3-8B Yang et al. (2025). These models were fine-tuned using the same methodology and dataset as the main PsyLLM model. As shown in Fig. 7 (Left), there is a clear trend of improving performance with increasing model size. The Qwen3-8B model achieves the highest overall Normalized Average score (0.757), indicating superior performance compared to the smaller models. The Qwen3-3B model performs better than the 1.7B version, achieving a Normalized Average of 0.738, while the smallest Qwen3-1.7B consistently shows the lowest overall performance with a Normalized Average of 0.675. These results suggest that larger base models, even with identical fine-tuning, exhibit stronger capabilities for generating nuanced and authentic psychological counseling responses.
@@
To investigate the influence of training data volume, we trained PsyLLM using varying proportions (20%, 40%, 60%, 80%, and 100%) of the OpenR1-Psy dataset, keeping the base model (Qwen3-8B) consistent. As shown in Fig. 7 (Right), the impact of data scale on the Normalized Average is not strictly monotonic. Performance generally improves from 20% (0.718) up to 80% (0.746), with 80% achieving the highest Normalized Average. However, the Normalized Average drops significantly when using the full 100% of the training data (0.656). The Safety metric remained consistently high across all data scales (around 0.99). This indicates that simply using more data for model training does not always yield the best performance.
@@
4.5 Case study
In this test case, a client reported persistent feelings of being under scrutiny, expressing the belief that as a Black woman, there is “no room for mistakes”. This belief contributed to ongoing exhaustion and emotional strain.
We compare the responses of CPsyCounX and our PsyLLM model. As shown in Figure 8, CPsyCounX primarily offers generic coping suggestions, such as seeking professional support or engaging in relaxation activities. While such strategies may provide momentary relief, they rely heavily on surface-level empathy and external recommendations, without addressing the client’s underlying cognitive and emotional patterns.
By contrast, PsyLLM demonstrates deeper clinical reasoning. Guided by DSM/ICD-informed understanding, the model identifies the client’s internalized perfectionism and hypervigilance as maladaptive thought patterns. It then integrates elements from cognitive behavioral therapy (CBT) to reframe rigid beliefs, while also drawing on humanistic principles of self-compassion to validate the client’s lived experience. Furthermore, the response invites reflection on rest and self-worth beyond external expectations, creating a supportive therapeutic space.
This example highlights PsyLLM’s capacity to move beyond mechanical empathy by combining diagnostic reasoning with multi-modal therapeutic strategies. Such integration aligns the model more closely with real-world clinical counseling standards, ensuring that support is not only empathetic but also clinically grounded.
@@ 5 Conclusion
This paper addresses the critical limitations of existing LLMs for mental health support, specifically, their lack of explicit diagnostic reasoning (DSM/ICD aligned) and diverse therapeutic modality integration. We introduce PsyLLM, the first model designed to systematically integrate both, facilitated by a novel automated data synthesis pipeline creating the OpenR1-Psy dataset. This dataset incorporates explicit reasoning traces guided by diagnostic standards and multiple therapeutic frameworks (e.g., CBT, ACT, psychodynamic), validated through multi-dimensional filtering. Our benchmark, assessing counseling quality, demonstrates that PsyLLM significantly outperforms state-of-the-art baselines.
Appendix A Related Work
A.1 Mental Support Chatbots
A.2 Deep Reasoning in LLMs
Deep reasoning in LLMs has attracted increasing attention, particularly since the introduction of CoT prompting Wei et al. (2022), which encourages models to generate intermediate reasoning steps before producing final answers. This technique has shown notable improvements on arithmetic and symbolic tasks, revealing the latent ability of LLMs to perform multi-step inference. Recent work has extended deep reasoning capabilities beyond symbolic computation into more complex domains such as math reasoning, medical understanding, multimodal reasoning, and emotional understanding Zhou et al. (2025b, a); Hu et al. (2025); Zhou et al. (2024). For example, Dou et al. (2025) proposes a dual-system framework inspired by Kahneman’s theory of fast and slow thinking, integrating reinforcement learning and vision-language reasoning to support adaptive decision-making. Similarly, reinforcement learning has been leveraged to enhance both text-based reasoning and multimodal emotion recognition, with modality-specific contributions disentangled through verifiable rewards Zhao et al. (2025). In mathematical contexts, Wang et al. (2025) introduces a model that bridges vision and code by aligning image understanding with code generation, achieving superior performance in structured geometry problems compared to models like GPT-4o. As research pushes toward more interactive and human-centric applications, deep reasoning in dialogue systems has emerged as a challenging yet crucial frontier. Maintaining coherence, adaptability, and emotional sensitivity across multiple conversational turns introduces new demands. To address this, Shu et al. (2025) reformulates CoT prompting as a dialogue between reasoning agents and applies reinforcement learning to improve both coherence and solution diversity. In the context of empathetic conversations, Xie et al. (2025b) proposes an LPE framework that combines CoT with emotional perception from spoken input, thereby guiding emotionally intelligent response generation. Moreover, psychological structures have been introduced to support deep reasoning in personality-related tasks: Yang et al. (2023) models psychological questionnaires as structured CoT chains for personality trait detection through multi-turn dialogue, while Chen et al. (2025) introduces a graph-based reasoning visualization system to facilitate collaboration between AI and therapists in therapeutic settings.
Appendix B Extended Multi-Turn Dialogue Case Study
To provide a more comprehensive illustration of PsyLLM’s capabilities in a multi-turn dialogue context, this section presents an extended example in Figure 9. This case study demonstrates how the model’s reasoning and responses evolve over several turns, maintaining contextual coherence and adapting its therapeutic strategy. Each counselor utterance is preceded by the model’s explicit reasoning trace, offering a transparent view into the clinical thought process that guides the dialogue, from initial validation to deeper exploration of the user’s feelings and patterns.
Appendix C Details for Human evaluator
We engaged five psychology experts to conduct human evaluation. Compensation was determined by the type and volume of work completed, in line with local regulations and professional standards.
Correlation Between Expert and Automatic Evaluations
To examine the consistency between human and automatic assessments, we randomly sampled a total of 500 test cases, evenly distributed across the four evaluation dimensions, from both our proposed model and baseline systems. For each dimension, we computed Spearman’s rank correlation between scores assigned by psychology experts and those generated by Gemini-2.0-Flash. As reported in Table 6, all correlations were statistically significant, confirming that automatic evaluation provides a reliable approximation of expert judgments and thereby supports the robustness of our evaluation framework.
| Metric | Correlation |
| Empathy & Insight | 0.585945∗∗ |
| Support & Autonomy | 0.534870∗∗ |
| Attunement & Presence | 0.566259∗∗ |
| Safety & Boundaries | 0.440841∗∗ |
| OA | 0.604751∗∗ |
Appendix D Detailed Automatic Evaluation Results
We report the complete automatic evaluation results of PsyLLM and baseline models across different topics in Table 7.
Appendix E Ethics Statements
@@
Data Availability Statement
The dataset introduced in this study, OpenR1-Psy, will be publicly released upon the acceptance of this paper. It will be hosted in an open-access repository and made freely available for non-commercial academic research purposes.
Appendix F Prompts for OpenR1-Psy
The OpenR1-Psy dataset employs a series of structured prompts for data construction and quality control. Figure 10 presents the prompt for Therapeutic Interaction Planning, Figure 11 for Empathic Dialogue Reconstruction, Figure 12 for ICD/DSM-Grounded Therapeutic Empathy, Figure 13 for Therapy-Guided Dialogue Validation, and Figure 14 for Therapy-Oriented Diagnostic Classification.
Appendix G Automatic Evaluation
For automatic evaluation, PsyLLM adopts several assessment prompts covering dialogue quality and reasoning evaluation. Figure 15 shows the prompt for Data Evaluation, Figure 16 for Performance Evaluation, and Figure 17 for Internal Thinking Evaluation. The evaluation standard refers to the evaluation metrics shows in Table 12.
Appendix H Model Comparison Examples
We provide qualitative case studies to compare the conversational behaviors of different psychological dialogue models. Figure 18 shows a dialogue between a client and DeepSeek-V3, Figure 19 with GPT-4o, Figure 20 with CPsyCounX, Figure 21 with ChatCounselor, Figure 22 with MeChat, and Figure 23 with PsyDTLLM.
| Topic | Model | Emp&In | Sup&Aut | Att&Pre | Saf&Bou |
| Self-growth | PsyDTLLM | 0.172 | 1.210 | 0.726 | 0.936 |
| MeChat | 0.013 | 0.866 | 0.357 | 0.904 | |
| ChatCounselor | 0.000 | 0.774 | 0.226 | 0.419 | |
@@ Appendix H Model Comparison Examples
| Emotion & Stress | PsyDTLLM | 0.213 | 1.286 | 0.801 | 0.875 |
| MeChat | 0.025 | 0.869 | 0.425 | 0.796 | |
| ChatCounselor | 0.016 | 0.873 | 0.337 | 0.504 | |
@@
| Education | PsyDTLLM | 0.105 | 1.012 | 0.581 | 0.802 |
| MeChat | 0.023 | 0.837 | 0.395 | 0.779 | |
| ChatCounselor | 0.000 | 0.794 | 0.305 | 0.412 | |
@@
| Love & Marriage | PsyDTLLM | 0.111 | 1.031 | 0.646 | 0.850 |
| MeChat | 0.016 | 0.833 | 0.440 | 0.796 | |
| ChatCounselor | 0.018 | 0.788 | 0.343 | 0.472 | |
@@
| Family Relationship | PsyDTLLM | 0.130 | 1.097 | 0.629 | 0.873 |
| MeChat | 0.017 | 0.870 | 0.407 | 0.770 | |
| ChatCounselor | 0.021 | 0.785 | 0.343 | 0.453 | |
@@
| Social Relationship | PsyDTLLM | 0.157 | 1.206 | 0.725 | 0.912 |
| MeChat | 0.010 | 0.814 | 0.324 | 0.804 | |
| ChatCounselor | 0.017 | 0.922 | 0.353 | 0.457 | |
@@
| Career | PsyDTLLM | 0.154 | 1.169 | 0.769 | 0.908 |
| MeChat | 0.015 | 1.031 | 0.523 | 0.846 | |
| ChatCounselor | 0.000 | 0.837 | 0.279 | 0.395 | |
@@
| Emp&In | Sup&Aut | Att&Pre | Saf&Bou | Normalized Avg | |
@@
| Emp&In | Sup&Aut | Att&Pre | Saf&Bou | Normalized Avg | |
@@
| Method | Metrics | ||||
| Emp&In | Sup&Aut | Att&Pre | Saf&Bou | Normalized Avg | |
| Prompt w/o Clinical Frame | 1.075 | 2.994 | 2.079 | 0.979 | 0.740 |
| Prompt w/o Therapy Guidance | 0.803 | 2.871 | 2.053 | 0.985 | 0.697 |
| OpenR1-Psy | 1.083 | 3.277 | 2.299 | 0.988 | 0.778 |
| Method | Emp&In | Sup&Aut | Att&Pre | Saf&Bou |
| In-Context Learning | 0.244 | 1.000 | 0.268 | 0.649 |
| Two-Phase Prompt | 0.592 | 1.450 | 0.669 | 0.698 |
| Ours | 1.059 | 2.787 | 2.012 | 0.988 |
| Perspective | Criterion | Score | Subtotal | References |
| Empathy & Insight | 1.1 Accurate Empathy for Core Emotions: Does the response accurately identify and validate the client’s specific core emotions (e.g., anxiety, powerlessness, self-doubt) with precise language, going beyond generic phrases? | 1 | ||
| 1.2 Insight into Underlying Needs & Pain Points: Does the response show insight into the client’s deeper, often unstated, emotional pain points or unmet needs (e.g., for safety, belonging, understanding) and gently guide exploration or affirm their validity? | 1 | 2 | Rogers (1957) | |
| Support & Autonomy | 2.1 Emotional Attunement & Validation: Does the counselor deeply empathize with and validate the client’s emotional experience, making the client feel "felt" and understood? | 1 | ||
| 2.2 Natural Integration of Support: Is supportive guidance (reflection, affirmation, gentle questions) woven naturally into the conversation, without jargon or explicit theory-naming? | 1 | |||
| 2.3 Respectful & Non-Directive Style: Is the language consistently warm, respectful, and non-judgmental, avoiding commands or prescriptive "you should" statements? | 1 | 4 | Martin (2013) | |
| 2.4 Gentle Guidance & Client Agency: Are suggestions offered tentatively, encouraging client-led exploration and small steps, rather than providing direct solutions and respecting client autonomy? | 1 | |||
| Attunement & Presence | 3.1 Conversational Fluency: Does the counselor communicate in a fluent, natural way, mirroring real therapeutic dialogue? | 1 | ||
| 3.2 Relational Presence: Is there evidence of genuine emotional connection and relational presence with the client? | 1 | 3 | Bordin (1979) | |
| 3.3 Pacing and Responsiveness: Does the counselor maintain appropriate pacing, leaving space for client expression, and respond in a natural, conversational manner to the client’s immediate input? | 1 | |||
| Safety & Boundaries | 4.1 Ethical Boundaries & Client Protection: Does the counselor avoid intrusive or leading language, respect privacy by not probing for unnecessary sensitive details, and prioritize client well-being and autonomy? | 0.5 | ||
| 4.2 Fostering Psychological Safety: Does the counselor consistently convey unconditional respect and acceptance, creating an environment free of judgment or dismissal, where the client feels safe to be authentic? | 0.5 | 1 | Guo et al. (2024) |